OpenStudy (anonymous):

A system of equations is shown below. -3x + 7y = -16 -9x + 5y = 16 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. Part B: Show that the equivalent system has the same solution as the original system of equations.

OpenStudy (anonymous):

let: -3x + 7y = -16 (1) -9x + 5y = 16 (2) For Part A: You can just multiply the first equation by -3.

OpenStudy (anonymous):

Then add the result to the second equation

OpenStudy (anonymous):

solve the system for y.

OpenStudy (anonymous):

Then take the value of y that you got and substitute it into either of the original two equations; then solve for x.